OpenStudy (anonymous):

Indicate which property is illustrated in Step 4. Step 1 12 - 5x - 12 + 6x = (12 - 5x) + (-12 + 6x) Step 2 = (12 - 5x) + (6x - 12) Step 3 = 12 + (-5x + 6x) - 12 Step 4 = 12 + (-5 + 6)x - 12 Step 5 = 12 + x - 12 Step 6 = x + 12 - 12 Step 7 = x + 0 Step 8 = x distributive associative identity commutative

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@croslali

OpenStudy (anonymous):

@SACAPUNTAS

OpenStudy (anonymous):

We did this one already. It should look familiar?

OpenStudy (anonymous):

comm again @SACAPUNTAS

OpenStudy (anonymous):

I'm here!

OpenStudy (anonymous):

i was asking if it commutative

OpenStudy (anonymous):

Oh! No. Distributive. That's the one that lets you say: \(ac + bc = (a + b)c\)

OpenStudy (anonymous):

Indicate which property is illustrated in Step 4. Step 1 12 + 11x - 12 - 2x + 5 = 12 + (11x - 12) - 2x + 5 Step 2 = 12 + (-12 + 11x) - 2x + 5 Step 3 = [12 + (-12)] + (11x - 2x) + 5 Step 4 = 0 + (11x - 2x) + 5 Step 5 = (11x - 2x) + 5 Step 6 = (11 - 2)x + 5 Step 7 = 9x + 5 associative additive identity distributive additive inverse

OpenStudy (anonymous):

Ok, do you know what property says \(a + -a = 0\)?

OpenStudy (anonymous):

ummmmmm inverse

OpenStudy (anonymous):

Yes, got it.

OpenStudy (anonymous):

-8(8m)=(-8 x 8 )mIndicate which property is illustrated below. commutative identity associative distributive

OpenStudy (anonymous):

@SACAPUNTAS

OpenStudy (anonymous):

I have no idea what that's even trying to show....

OpenStudy (anonymous):

lol Indicate which property is illustrated below. -8(8m)=(-8 x 8 )m commutative identity associative distributive

OpenStudy (anonymous):

@SACAPUNTAS

OpenStudy (anonymous):

Uh... that's just...

OpenStudy (anonymous):

Lemme think a sec.

OpenStudy (anonymous):

Oh. Yeah. Associative, I think.